### v1

# Adaptive Task Space Non-Singular Terminal Super-Twisting Sliding Mode Control of a 7-DOF Robotic Manipulator

Abstract
This paper presents a new task-space Non-singular Terminal Super-Twisting Sliding Mode (NT-STSM) controller with adaptive gains for robust trajectory tracking of a 7-DOF robotic manipulator. The proposed approach addresses the challenges of chattering, unknown disturbances, and rotational motion tracking, making it suited for high-DOF manipulators in dexterous manipulation tasks. A rigorous boundedness proof is provided, offering gain selection guidelines for practical implementation. Simulations and hardware experiments with external disturbances demonstrate the proposed controller’s robust, accurate tracking with reduced control effort under unknown disturbances compared to other NT-STSM and conventional controllers. The results demonstrated that the proposed NT-STSM controller mitigates chattering and instability in complex motions, making it a viable solution for dexterous robotic manipulations and various industrial applications.
Index Terms:
Adaptive control, non-singular terminal sliding mode control, 7-DOF robotic manipulator, super-twisting sliding mode control, task space control.
I Introduction
The development of robust control algorithms is necessary for industrial robotic manipulators in applications such as remote surgery, cooperative multi-robot manipulation, and handling varying payloads. These applications require precise trajectory tracking, robustness to disturbances, and energy-efficient control strategies. High degree-of-freedom (DOF) manipulators offer an extensive range of motion, however, their complex nonlinear dynamics, with model uncertainties and external disturbances, pose significant control challenges.
Sliding mode (SM) control provides precise tracking capabilities and is robust against disturbances and uncertainties. Non-singular Terminal Sliding Mode (NTSM) control offers fast, finite-time convergence speed and avoids singularities, making it appealing for robotic manipulators [1]. However, its inherent chattering effect, caused by high-frequency switching, limits its practical implementation. Super-Twisting Sliding Mode (STSM) control [2] alleviates this issue by smoothing the control signal, making it suitable for robotic applications [3]. Adaptive gain STSM controllers have been developed to further mitigate chattering, improve tracking performance, and alleviate the requirement of prior knowledge of the disturbance bounds [4, 5]. Combining the NTSM and STSM control into a Non-singular Terminal Super-Twisting Sliding Mode (NT-STSM) controller has been explored in [6, 7, 8, 9]. However, existing studies are largely limited to low-DOF manipulators and simulations without extensive experimental verifications, indicating a gap in practical development and real-world validation for high-DOF robotic manipulators.
Task-space control, where the reference trajectory is defined in Cartesian space, is important for industrial robotic tasks such as precision assembly, force-controller interactions, and collaborative manipulation [10]. Unlike joint-space control, which requires inverse kinematics and dynamic transformations, task-space controllers directly regulate the end-effector motion, improving accuracy and ease of implementation in practical settings. However, designing a robust task-space SM control strategy that effectively handles both translational and rotational motion remains challenging. While SM joint space controllers are common [11, 12], they require converting task space objectives into joint space and are prone to modeling errors and uncertainties [13]. Prior task-space SM control approaches are limited to low-DOF manipulators [13] or focus only on the translational motion [14, 15], failing to address the increased chattering risk in the rotational motion.
This paper introduces a task-space NT-STSM controller for high-DOF robotic manipulators, addressing the limitations of prior task-space SM control methods and practical NT-STSM implementations. In summary, the proposed approach applies to the control of robotic manipulators with any DOFs for dexterous manipulations. The main contributions are:
  1. 1.
    An NT-STSM controller is proposed with an adaptive gain law designed to improve tracking performance and reduce chattering without requiring prior knowledge of disturbance bounds. A novel boundedness proof is derived with tunable gain selection guidelines for practical implementation. Compared to [9], the proposed boundedness proof relaxes the relationship between the STSM gains, providing improved performance on a 7-DOF robotic manipulator.
  2. 2.
    The proposed NT-STSM controller is designed in task space with unit quaternions representing orientation. Unlike prior task-space SM control approaches [13, 14, 15], this work explicitly incorporates unit quaternions, allowing for rotational motion while preventing kinematic singularities. As demonstrated in Section IV, the rotational motion can be more prone to chattering, reinforcing the importance of the proposed chatter-free task space NT-STSM controller.
  3. 3.
    The proposed NT-STSM controller’s tracking performance and robustness against unknown disturbances are validated with simulations and experiments on a 7-DOF Franka Emika Robot with extensive comparisons. The proposed controller demonstrates lower position and orientation tracking error compared to conventional PD control with similar control effort and the NT-STSM controller developed in [9]. Unlike NTSM and STSM controllers, which suffer from severe chattering and instability, the proposed controller maintains smooth trajectory tracking.
II Preliminaries
II-A Dynamic Model
Consider a $7$-DOF robot manipulator with the joint space dynamics of the form [16, Ch. 3]
$\mathcal{M}(\bm{q})\ddot{\bm{q}}+\mathcal{C}(\bm{q},\dot{\bm{q}})\dot{\bm{q}}+\mathcal{G}(\bm{q})=\bm{\tau}+\bm{\tau}_{d}(\bm{q}),$ (1)
where $\bm{q},\dot{\bm{q}},\ddot{\bm{q}}\in\mathbb{R}^{7}$ are the joint position, velocity, and acceleration vectors, respectively, $\mathcal{M}(\bm{q})\in\mathbb{R}^{7\times 7}$ is the inertia matrix, $\mathcal{C}(\bm{q},\dot{\bm{q}})\in\mathbb{R}^{7\times 7}$ is the Coriolis and centripetal torque matrix, $\mathcal{G}(\bm{q})\in\mathbb{R}^{7}$ is the gravitational torque vector, $\bm{\tau}\in\mathbb{R}^{7}$ is the joint torque control input vector, $\bm{\tau}_{d}(\bm{q})=\bm{\tau}_{f}(\bm{q})+\bm{\tau}_{u}(\bm{q},\dot{\bm{q}})+\bm{\tau}_{e}\in\mathbb{R}^{7}$ is the summation of the internal torque friction, $\bm{\tau}_{f}(\bm{q})$, the unmodelled dynamics due to modelling inaccuracies, $\bm{\tau}_{u}(\bm{q},\dot{\bm{q}})$, and the external torque exerted on the manipulator, $\bm{\tau}_{e}$. In task space, the dynamics of the robot manipulators may be represented as [16, Ch. 3]
¯$\bar{\mathcal{M}}(\bm{q})\ddot{\bm{x}}+\bar{\mathcal{C}}(\bm{q},\dot{\bm{q}})+\bar{\mathcal{G}}(\bm{q})=\bm{u}+\bm{f}_{d}(\bm{q}).$ (2)
The end-effector pose in task space is denoted as $\bm{x}=\begin{bmatrix}\bm{p}^{T},&\bm{\xi}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{7}$ where $\bm{p}\in\mathbb{R}^{3}$ is the translational position and $\bm{\xi}=\begin{bmatrix}\eta,&\bm{\epsilon}^{T}\end{bmatrix}^{T}\in\mathrm{S}^{3}\subset\mathbb{R}^{4}$ is the orientation in the form of a unit quaternion. The quaternion is made up of $\eta$, a scalar denoting the real part, and $\bm{\epsilon}=\begin{bmatrix}\epsilon_{x},&\epsilon_{y},&\epsilon_{z}\end{bmatrix}^{T}$, a vector denoting the imaginary part. The end-effector velocity and acceleration are denoted as $\dot{\bm{x}}=\begin{bmatrix}\dot{\bm{p}}^{T},&\bm{\omega}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{6}$ and $\ddot{\bm{x}}=\begin{bmatrix}\ddot{\bm{p}}^{T},&\dot{\bm{\omega}}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{6}$, where $\bm{\omega},\ \dot{\bm{\omega}}\in\mathbb{R}^{3}$ are the angular velocity and acceleration of the end-effector, respectively.
The end-effector velocity and joint velocity are related by the Jacobian matrix, as $\dot{\bm{x}}=J(\bm{q})\dot{\bm{q}}$. For the remainder of the paper, the dependencies of the dynamic terms on $\bm{q}$ and $\dot{\bm{q}}$ will be omitted.
Property 1
The Cartesian parametric matrices in (2) are related to the joint space matrices in (1) by [16, Ch. 3]
$\bm{\tau}$ $=J^{T}\bm{u},$$\bar{\mathcal{M}}=(J\mathcal{M}^{-1}J^{T})^{-1},$
$\bar{\mathcal{C}}$ =¯$=\bar{\mathcal{M}}(J\mathcal{M}^{-1}\mathcal{C}\dot{\bm{q}}-\dot{J}\dot{\bm{q}}),$ $\bar{\mathcal{G}}=\bar{\mathcal{M}}J\mathcal{M}^{-1}\mathcal{G}.$
Property 2
The inertia matrix, $\mathcal{M}$, is positive-definite symmetric, ˙$\dot{\mathcal{M}}-2\mathcal{C}$ is skew-symmetric, and the Jacobian matrix, $J$, has full rank [16, Ch. 3].
II-B Trajectory Generation
The trajectory generator provides the controller with a desired trajectory, $\begin{bmatrix}\bm{x}_{d},&\dot{\bm{x}}_{d},&\ddot{\bm{x}}_{d}\end{bmatrix}^{T}$. This trajectory is interpolated between an initial pose, $\bm{x}_{d_{0}}$, and goal pose, $\bm{x}_{d_{g}}$.
The translational motion is calculated using the clamped cubic spline approach to provide smooth trajectories with zero velocities at the start and end. Unit quaternions represent the orientation position, therefore, standard interpolation methods may not result in consistent and smooth rotational trajectories [17]. To interpolate between an initial quaternion, $\bm{\xi}_{d_{0}}$, and a goal quaternion $\bm{\xi}_{d_{g}}$, the cubic Hermite curve interpolation method [18] is used:
$\bm{\xi}_{d}(t)$ $=\frac{F_{a2q}(WF_{q2a}(\bm{\xi}_{d_{g}}\times\bm{\xi}_{d_{0}}^{-1})\times\bm{\xi}_{d_{0}})}{|F_{a2q}(WF_{q2a}(\bm{\xi}_{d_{g}}\times\bm{\xi}_{d_{0}}^{-1})\times\bm{\xi}_{d_{0}})|},$ (3)
$W$ $=3\left(\frac{t}{T}\right)^{2}-2\left(\frac{t}{T}\right)^{3},$
where $T$ is the time between two poses, $\bm{\xi}_{1}\times\bm{\xi}_{2}$ represents quaternion multiplication, $|\bm{\xi}|$ represents the norm of the quaternion, and $\bm{\xi}^{-1}$ represents the inverse of the quaternion, which in the case of unit quaternions is equal to the conjugate of the quaternion, $\bm{\xi}^{*}$. $F_{q2a}(\bm{\xi})$ represents a function to convert a quaternion to a scaled angle axis, $\bm{\Theta}$, and $F_{a2q}(\bm{\Theta})$ represents a function to convert a scaled angle axis to a quaternion.
The desired angular velocity is calculated from (3) using the quaternions as [19]
$\bm{\omega}_{d}(t)=\frac{2}{\Delta t}\begin{bmatrix}\eta^{\prime}\epsilon_{x}-\epsilon_{x}^{\prime}\eta-\epsilon_{y}^{\prime}\epsilon_{z}+\epsilon_{z}^{\prime}\epsilon_{y}\\ -\eta^{\prime}\epsilon_{y}-\epsilon_{x}^{\prime}\epsilon_{z}+\epsilon_{y}^{\prime}\eta+\epsilon_{z}^{\prime}\epsilon_{x}\\ -\eta^{\prime}\epsilon_{z}+\epsilon_{x}^{\prime}\epsilon_{y}-\epsilon_{y}^{\prime}\epsilon_{x}+\epsilon_{z}^{\prime}\eta\end{bmatrix},$ (4)
where $\Delta t$ is the sampling time, $\eta=\eta_{d}(t)$, $\bm{\epsilon}=\bm{\epsilon}_{d}(t)$, $\eta^{\prime}=\eta_{d}(t-\Delta t)$, and $\bm{\epsilon}^{\prime}=\bm{\epsilon}_{d}(t-\Delta t)$. The desired angular acceleration is calculated using the finite difference method.
II-C Third Order Sliding Mode Observer
SM observers provide a robust method of estimating the velocity of robotic manipulators. According to experimental tests in [20], SM observers exhibit consistent and high accuracy over a range of sensor resolutions and sampling rates, whereas moving average and derivative filters often had significant estimation errors. SM observers have the additional benefit of being robust to inconsistent sensor readings [20]. Similar to SM controllers, SM observers suffer from chattering in real-world applications. An approach to reducing chatter is to use a TOSM observer, although they have slower convergence than lower-order observers. In [21], linear terms are introduced to increase the convergence speed. By assigning $\bm{g}_{1}=\bm{x}$, $\bm{g}_{2}=\dot{\bm{x}}$, and denoting $\hat{\bm{g}}_{1}$ and $\hat{\bm{g}}_{2}$ as the estimated states, a TOSM observer is adapted from [21] and formulated in task space
$\dot{\hat{\bm{g}}}_{1}=$ $\ \hat{\bm{g}}_{2}+\alpha_{o2}\odot|\bm{g}_{1}-\hat{\bm{g}}_{1}|^{\frac{2}{3}}\odot\textrm{sign}(\bm{g}_{1}-\hat{\bm{g}}_{1})$
$+k_{o2}(\bm{g}_{1}-\hat{\bm{g}}_{1}),$ (5)
$\dot{\hat{\bm{g}}}_{2}=$$\ \bar{\mathcal{M}}^{-1}(\bm{u}-\bar{\mathcal{C}}-\bar{\mathcal{G}})+\alpha_{o1}\odot|\dot{\hat{\bm{g}}}_{1}-\hat{\bm{g}}_{2}|^{\frac{1}{2}}\odot\textrm{sign}(\dot{\hat{\bm{g}}}_{1}-\hat{\bm{g}}_{2})$
$+k_{o1}(\bm{g}_{1}-\hat{\bm{g}}_{1})+\hat{\bm{z}},$ (6)
$\dot{\hat{\bm{z}}}=$ $\ \alpha_{o0}\textrm{sign}(\dot{\hat{\bm{g}}}_{1}-\hat{\bm{g}}_{2}),$ (7)
where $\alpha_{o2}$, $\alpha_{o1}$, $\alpha_{o0}\in\mathbb{R}^{7\times 1}$ are positive gain vectors, $k_{o2}$, $k_{o1}$ are positive scalar gains, and $\odot$ is the Hadamard product. The estimation error of the observed states are defined as $\tilde{\bm{g}}_{1}=\bm{g}_{1}-\hat{\bm{g}}_{1}$, and $\tilde{\bm{g}}_{2}=\bm{g}_{2}-\hat{\bm{g}}_{2}$, detailed in [21, Eq. (29)]. The proof of convergence for the TOSM observer is in [21, Theorem 1], demonstrating that $\tilde{\bm{g}}_{1}$ and $\tilde{\bm{g}}_{2}$ are stable and converge to zero in finite time. In implementation, first-order approximation methods are used to integrate the observer terms.
Despite introducing the linear terms to the observer to increase the convergence speed, there is still an initial period during which the estimated states are inaccurate before they converge. This can lead to unpredictable behaviour of the manipulator. A finite difference method and exponential moving average filter are used to estimate the velocity during this convergence period as
$\dot{\bar{\bm{x}}}(t)=\alpha_{e}\frac{\bm{x}(t)-\bm{x}(t-\Delta t)}{\Delta t}+(1-\alpha_{e})\dot{\bar{\bm{x}}}(t-\Delta t),$
where $\alpha_{e}$ is a positive filter design parameter. The position and velocity sent to the controller are selected as
$\hat{\bm{x}}$ =$=\begin{cases}\bm{x}&\text{if }|\hat{\bm{g}}_{1}-\bm{x}|>\eta_{q}\\ \hat{\bm{g}}_{1}&\text{if }|\hat{\bm{g}}_{1}-\bm{x}|\leq\eta_{q}\end{cases},$ (8)
$\dot{\hat{\bm{x}}}$ =$=\begin{cases}\dot{\bar{\bm{x}}}&\text{if }|\hat{\bm{g}}_{2}-\dot{\bar{\bm{x}}}|>\eta_{q}\\ \hat{\bm{g}}_{2}&\text{if }|\hat{\bm{g}}_{2}-\dot{\bar{\bm{x}}}|\leq\eta_{q}\end{cases},$ (9)
where $\eta_{q}>0$ is a positive design parameter.
III Adaptive Robust Controller Design
In this section, a novel NT-STSM controller is proposed along with rigorous boundedness proof. The block diagram of the controller and observer is shown in Fig. 1.
Figure 1: Block diagram of the proposed control design.
III-A Non-Singular Terminal Super-Twisting Sliding Mode Controller
Compared to the conventional SM control, an NTSM controller offers appealing characteristics of fast, finite time convergence and avoiding singularities [1]. Consider the nonlinear sliding surface as
$\bm{s}=\bm{e}+\beta\dot{\bm{e}}^{\alpha},$ (10)
where $\beta>0$, $\alpha=\iota_{1}/\iota_{2}$, and $\iota_{1}>0$, $\iota_{2}>0$ are adjacent odd numbers such that $21<\alpha<$ 2, and
$\bm{e}=\begin{bmatrix}\tilde{\bm{p}}\\ \tilde{\bm{\epsilon}}\end{bmatrix},\quad\dot{\bm{e}}=\begin{bmatrix}\dot{\tilde{\bm{p}}}\\ \dot{\tilde{\bm{\epsilon}}}\end{bmatrix}=\begin{bmatrix}\dot{\tilde{\bm{p}}}\\ \left(\tilde{\eta}I+[\tilde{\bm{\epsilon}}]\right)\tilde{\bm{\omega}}\end{bmatrix}=H\begin{bmatrix}\dot{\tilde{\bm{p}}}\\ \tilde{\bm{\omega}}\end{bmatrix},$ (11)
where
$\tilde{\bm{p}}$ $=\hat{\bm{p}}-\bm{p}_{d},$ $\dot{\tilde{\bm{p}}}$ $=\dot{\hat{\bm{p}}}-\dot{\bm{p}}_{d},$
$\tilde{\eta}$ $=\hat{\eta}\eta_{d}+\hat{\bm{\epsilon}}^{T}\bm{\epsilon}_{d},$ $\tilde{\bm{\epsilon}}$ $=-\hat{\eta}\bm{\epsilon}_{d}+\eta_{d}\hat{\bm{\epsilon}}-[\hat{\bm{\epsilon}}]\bm{\epsilon}_{d},$
$\tilde{\bm{\omega}}$ $=\hat{\bm{\omega}}-\bm{\omega}_{d},$ $H$ $=\begin{bmatrix}I&0\\ 0&\left(\tilde{\eta}I+[\tilde{\bm{\epsilon}}]\right)\end{bmatrix},$
and $[\bm{\epsilon}]=\begin{bmatrix}0&-\epsilon_{z}&\epsilon_{y}\\ \epsilon_{z}&0&-\epsilon_{x}\\ -\epsilon_{y}&\epsilon_{x}&0\end{bmatrix}$ is the skew-symmetric matrix operator. The formulation of the quaternion error, $\tilde{\bm{\epsilon}}$ and $\dot{\tilde{\bm{\epsilon}}}$ can be found in [22, Ch. 3].
The main challenge in using a conventional SM controller in real-world applications is the chattering effect. A second-order SM controller is a method of reducing chattering without reducing the controller’s performance. The STSM controller is a second-order method that introduces a continuous function of the sliding surface and an integral of the discontinuous function of the sliding surface to minimize chattering. The controller is designed as
$\bm{u}=$$\ \bar{\mathcal{M}}\left(\ddot{\bm{x}}_{d}+H^{-1}\left(\frac{-\dot{\bm{e}}^{(2-\alpha)}}{\alpha\beta}-\kappa_{1}|S|^{\frac{1}{2}}\textrm{sign}(\bm{s})+\bm{\nu}\right)\right)$
$+\bar{\mathcal{C}}+\bar{\mathcal{G}},$ (12)
$\dot{\bm{\nu}}=$ $\ -\kappa_{2}\textrm{sign}(\bm{s}),$
where $|S|^{\frac{1}{2}}=\textrm{diag}(|\bm{s}|^{\frac{1}{2}})$ and
$\kappa_{1}=\begin{bmatrix}\kappa_{11}&0&0\\ 0&\ddots&0\\ 0&0&\kappa_{16}\end{bmatrix},\ \kappa_{2}=\begin{bmatrix}\kappa_{21}&0&0\\ 0&\ddots&0\\ 0&0&\kappa_{26}\end{bmatrix},$
are positive definite matrices. When implementing the controller on the manipulator, it is transformed into joint space, where $\bm{\tau}=J^{T}\bm{u}$.
Consider the invertibility of $H$. The determinant of $H$ is given by $\textrm{det}(H)=\tilde{\eta}(\tilde{\eta}^{2}+\tilde{\bm{\epsilon}}^{T}\tilde{\bm{\epsilon}})$. The error between two unit quaternions is derived from a quaternion product and must result in a unit quaternion [23]. According to the unit norm property of quaternions:
$\tilde{\eta}^{2}+\tilde{\bm{\epsilon}}^{T}\tilde{\bm{\epsilon}}=1,$
resulting in $\textrm{det}(H)=\tilde{\eta}$.
Assumption 1
It is assumed that the desired orientation trajectory is set to be sufficiently close to the current robot orientation. This is ensured by smooth interpolation in (3) and periodic checks of the robot’s current position by the trajectory generator. The proximity between the desired and actual orientations ensures that they do not result in orthogonal conditions or specific alignments that would cause $\textrm{det}(H)$ to be zero. Therefore, $\textrm{det}(H)=\tilde{\eta}\neq 0$, and $H$ remains invertible.
III-B Adaptive Law
Adaptive gains offer a dynamic response to uncertainties and disturbances, effectively compensating for them without requiring prior knowledge of their bounds [5]. This is important for systems characterized by uncertainties and disturbances that can vary over time. The adaptive law for the controller gains $\kappa_{1}$ and $\kappa_{2}$ in (12) is designed as
$\dot{L}_{i}$ =$=\begin{cases}-\eta_{\alpha},&\text{if }\kappa_{1i}\geq\kappa_{1\text{max}},\\ \omega_{\alpha}|s_{i}|\textrm{sign}(|s_{i}|-\mu_{\alpha}),&\text{if }\kappa_{1\text{max}}>\kappa_{1i}>\kappa_{1\text{min}},\\ \eta_{\alpha},&\text{if }\kappa_{1i}\leq\kappa_{1\text{min}},\end{cases}$ (13)
$\kappa_{1i}$ $=\Omega_{1}\sqrt{\frac{2\gamma}{(1-\theta)\Omega_{2}}}\sqrt{L_{i}},\qquad\kappa_{2i}=\frac{\theta+1}{1-\theta}L_{i},$ (14)
where $L\in\mathbb{R}^{6}$ is a positive adaptation parameter and $\omega_{\alpha}$, $\mu_{\alpha}$, $\eta_{\alpha}$, $\kappa_{1\text{max}}$, $\kappa_{1\text{min}}$, are constant positive adaptive law parameters and $\Omega_{1}$, $\Omega_{2}$, $\gamma$, and $\theta$ are constant positive scalar gain parameters that satisfy $10<\theta<$ 1 and
$\Omega_{1}\Gamma^{2}-2\frac{\Omega_{2}}{\gamma}\Gamma>\frac{1}{4}(\Gamma+\Omega_{1}\Gamma)^{2}-(1+\Omega_{1})\Omega_{2}\Gamma\theta+\Omega_{2}^{2},$ (15)
where $\Gamma=\alpha\beta\textrm{diag}(\dot{\bm{e}}^{(\alpha-1)})\in\mathbb{R}^{6}$. See Remarks 1 and 2 in Section III-C for tuning guidelines.
III-C Non-Singular Terminal Super-Twisting Sliding Mode Controller Convergence Analysis
This subsection presents the boundedness analysis considering the system (2) under the controller (12). First, the main theorem is presented. Then, the closed-loop dynamics are analyzed and reformulated such that the Lyapunov candidate function can be introduced. Next, the convergence of the sliding surface to a bounded region is analyzed. Finally, the tracking bounds of the position and velocity errors is analyzed when the sliding surface is zero and when it converges to a bounded region.
Lemma 1
[24] Assume a positive definite function, $V$, and its derivative satisfies
$\dot{V}\leq-\gamma_{1}V^{\gamma_{2}},$ (16)
where $\gamma_{1}>0$ and $10<\gamma_{2}<$ 1. The function $V$ converges to zero from any initial state within a finite time, determined by
$t_{f}(x_{0})\leq\frac{V^{1-\gamma_{2}}(x_{0})}{\gamma_{1}(1-\gamma_{2})}.$ (17)
Assumption 2
The lumped disturbances on the system, $\bm{\Pi}=H\bar{\mathcal{M}}^{-1}\bm{f}_{d}=HJ\mathcal{M}^{-1}\bm{\tau}_{d}$, are first-order differentiable and there exists a vector of positive constants, $\bm{\delta}\in\mathbb{R}^{6}$, such that $|\dot{\bm{\Pi}}|\leq\bm{\delta}$ is globally bounded.
Theorem 1
Consider the system dynamics (2) satisfying Assumption 2 with the control input (12). If the variable gains, $\kappa_{1}$ and $\kappa_{2}$ are selected as in (14), then in the presence of Lebesgue-measurable noise of maximal magnitude $\epsilon_{c}$, while selecting $\mu_{\alpha}>\epsilon_{c}$, a vicinity of the origin defined by $\bm{s}\leq\mu_{\alpha}$ is globally and robustly finite time stable for every value of perturbation derivative $|\dot{\bm{\Pi}}|\leq\bm{\delta}$. A trajectory starting at $\bm{s}(0)$ will converge to the region $|\bm{s}|\leq\mu_{\alpha}$ from any initial condition $|\bm{s}(0)|>\mu_{\alpha}$ in finite time. The upper bound on the time to reach this domain can be estimated as
$t_{f}(\bm{s}(0))\leq\frac{2}{\vartheta}V_{1}^{\frac{1}{2}}(\bm{s}(0)),$ (18)
where
$V_{1}=$ $|\bm{s}|-2\sqrt{\frac{(1-\theta)\Omega_{2}}{2\gamma L}}|\bm{s}|^{\frac{1}{2}}\textrm{sign}(\bm{s})(\bm{\nu}+\bm{\Pi})$
$+\frac{(1-\theta)\Omega_{2}}{2L}(\bm{\nu}+\bm{\Pi})^{2}+\frac{1}{2}(L-L^{*})^{2},$ (19)
and $\vartheta=\textrm{min}\{\chi,\sqrt{2}\omega_{\alpha}\mu_{\alpha}\}$, $\chi=\frac{\lambda^{\frac{1}{2}}_{\textrm{min}}(P)\lambda_{\textrm{min}}(\tilde{Q}_{R})}{\lambda_{\textrm{max}}(P)}$. Lastly, the finite time tracking control convergence is achieved in a finite time according to
$t_{f}(\bm{e})\leq\frac{2\alpha}{\beta^{-\frac{1}{\alpha}}(\alpha-1)}\frac{V_{3}(\bm{e}(0))^{\frac{\alpha-1}{2\alpha}}}{2^{\frac{1+\alpha}{2\alpha}}},$ (20)
where $V_{3}=\frac{1}{2}\bm{e}^{2}$. In the presence of measurement noise, the tracking errors converge to the following regions
$|{\bm{e}}|\leq 2\mu_{\alpha},\ \ |{\dot{\bm{e}}}|\leq\left(\frac{\mu_{\alpha}}{\beta}\right)^{\frac{1}{\alpha}}.$ (21)
Proof:
III-C1 Closed-loop Dynamics
This section introduces the closed-loop dynamics of the manipulator and arranges the system into an appealing form.
Given the system (2) with the controller (12), the closed-loop dynamics can be written as
$\ddot{\bm{x}}=\ddot{\bm{x}}_{d}+H^{-1}\left(\frac{-\dot{\bm{e}}^{(2-\alpha)}}{\alpha\beta}-\kappa_{1}|S|^{\frac{1}{2}}\textrm{sign}(\bm{s})+\bm{\nu}\right)+\bar{\mathcal{M}}^{-1}\bm{f}_{d}.$ (22)
The acceleration error is expressed as in [23]
$\ddot{\bm{e}}=\begin{bmatrix}\ddot{\tilde{\bm{p}}}\\ \ddot{\tilde{\bm{\epsilon}}}\end{bmatrix}=\begin{bmatrix}\ddot{\tilde{\bm{p}}}\\ \left(\tilde{\eta}I+[\tilde{\bm{\epsilon}}]\right)\dot{\tilde{\bm{\omega}}}\end{bmatrix},$ (23)
where $\dot{\tilde{\bm{\omega}}}=\dot{\bm{\omega}}-\dot{\bm{\omega}}_{d}$. Rearranging (22) results in
$\ddot{\bm{e}}$ $=H(\ddot{\bm{x}}-\ddot{\bm{x}}_{d})$
$=\frac{-\dot{\bm{e}}^{(2-\alpha)}}{\alpha\beta}-\kappa_{1}|S|^{\frac{1}{2}}\textrm{sign}(\bm{s})+\bm{\nu}+H\bar{\mathcal{M}}^{-1}\bm{f}_{d}.$ (24)
The derivative of the sliding surface (10) is derived as
$\dot{\bm{s}}$ $=\dot{\bm{e}}+\alpha\beta\textrm{diag}(\dot{\bm{e}}^{(\alpha-1)})\ddot{\bm{e}}$
$=\dot{\bm{e}}+\bm{\Gamma}\left(\frac{-\dot{\bm{e}}^{(2-\alpha)}}{\alpha\beta}-\kappa_{1}|S|^{\frac{1}{2}}\textrm{sign}(\bm{s})+\bm{\nu}+H\bar{\mathcal{M}}^{-1}\bm{f}_{d}\right)$
$=\bm{\Gamma}(-\kappa_{1}|S|^{\frac{1}{2}}\textrm{sign}(\bm{s})+\bm{\nu}+\bm{\Pi}).$ (25)
Due to the constraints on $\alpha$ and $\beta$ in (10), $\bm{\Gamma}=\alpha\beta\textrm{diag}(\dot{\bm{e}}^{(\alpha-1)})\geq 0$. Introducing new states as $\bm{y}_{1}=\bm{s}$ and $\bm{y}_{2}=\bm{\nu}+\bm{\Pi}$, we have
$\dot{\bm{y}}_{1}$ $=\dot{\bm{s}}=\Gamma(-\kappa_{1}|Y_{1}|^{\frac{1}{2}}\textrm{sign}(\bm{y}_{1})+\bm{y}_{2}),$ (26)
$\dot{\bm{y}}_{2}$ $=\dot{\bm{\nu}}+\dot{\bm{\Pi}}=-\kappa_{2}\textrm{sign}(\bm{y}_{1})+\dot{\bm{\Pi}},$ (27)
where $|Y_{1}|^{\frac{1}{2}}=\textrm{diag}(|\bm{y}_{1}|^{\frac{1}{2}})\in\mathbb{R}^{6\times 6}$.
Define $\bm{\zeta}$ as in [25], taking the $i$-th Cartesian dimension,
$\bm{\zeta}=\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T}=\begin{bmatrix}|y_{1i}|^{\frac{1}{2}}\textrm{sign}(y_{1i}),&y_{2i}\end{bmatrix}^{T}.$ (28)
For clarity, assume the $i$-th element for each term in the remainder of this section. Denoting $y_{1}\triangleq y_{1i},\ y_{2}\triangleq y_{2i},\ \kappa_{1}\triangleq\kappa_{1i},\ \kappa_{2}\triangleq\kappa_{2i},\ \Gamma\triangleq\Gamma_{i},\ \dot{\Pi}\triangleq\Pi_{i}$, then
$\dot{\zeta}_{1}$ $=\frac{1}{2}|y_{1}|^{-\frac{1}{2}}\dot{y}_{1}=\frac{1}{2}|y_{1}|^{-\frac{1}{2}}\Gamma(-\kappa_{1}|y_{1}|^{\frac{1}{2}}\textrm{sign}(y_{1})+y_{2})$
$=|y_{1}|^{-\frac{1}{2}}\frac{1}{2}\Gamma(-\kappa_{1}\zeta_{1}+\zeta_{2}),$ (29)
$\dot{\zeta}_{2}$ $=-\kappa_{2}\textrm{sign}(y_{1})+\dot{\Pi}=|y_{1}|^{-\frac{1}{2}}(-\kappa_{2}|y_{1}|^{\frac{1}{2}}\textrm{sign}(y_{1}))+\dot{\Pi}$
$=|y_{1}|^{-\frac{1}{2}}(-\kappa_{2}\zeta_{1})+\dot{\Pi}.$ (30)
Introducing a new variable $\pi(t,y)=\dot{\Pi}\textrm{sign}(y_{1})$, (29) and (30) can be arranged into
$\dot{\bm{\zeta}}=|y_{1}|^{-\frac{1}{2}}A\begin{bmatrix}\zeta_{1}\\ \zeta_{2}\end{bmatrix},$ (31)
where $A=\begin{bmatrix}-\frac{1}{2}\kappa_{1}\Gamma&\frac{1}{2}\Gamma\\ -(\kappa_{2}-\pi)&0\end{bmatrix}$, similar to the form in [2].
III-C2 Sliding Mode Dynamics Boundedness Analysis
The following analysis considers the boundedness of the sliding mode dynamics (31) while considering Lebesgue-measurable noise’s influence and adaptive gain selection (13).
Consider the following Lyapunov function:
$V_{1}$ $=V_{2}+\frac{1}{2}(L-L^{*})^{2},$ (32)
where $L^{*}$ is the upper bound of the gain values $L\leq L^{*}$, and
$V_{2}=\bm{\zeta}^{T}P\bm{\zeta}.$ (33)
The matrix $P$ is symmetric and positive definite as
$P=\begin{bmatrix}p_{11}&p_{12}\\ p_{21}&p_{22}\end{bmatrix}.$ (34)
As in [25], since $V_{2}$ is not locally Lipschitz, conventional versions of the Lyapunov theorem cannot be used here. However, $V_{2}$ is continuous along the state trajectories of $\bm{y_{1}}=\bm{s}$ and therefore the boundedness can be shown using Zubov’s Theorem [25, Appendix A]. Following Rayleigh’s Theorem as in [9], $V_{2}$ is positive definite and radially unbounded:
$\lambda_{\textrm{min}}(P)||\bm{\zeta}||_{2}^{2}\leq V_{2}\leq\lambda_{\textrm{max}}(P)||\bm{\zeta}||_{2}^{2},$ (35)
where $||\bm{\zeta}||_{2}^{2}=|y_{1}|+y^{2}_{2}$. Differentiating $V_{2}$ along the trajectories of the closed-loop dynamics (31)
$\dot{V}_{2}=-|y_{1}|^{-\frac{1}{2}}\bm{\zeta}^{T}\tilde{Q}\bm{\zeta}.$ (36)
Let $\tilde{Q}$ be defined as
$-\tilde{Q}=A^{T}P+PA,\ \ \tilde{Q}=\begin{bmatrix}\tilde{q}_{11}&\tilde{q}_{12}\\ \tilde{q}_{21}&\tilde{q}_{22}\end{bmatrix},$ (37)
which is computed as
${\tilde{Q}}:\begin{cases}\tilde{q}_{11}=\kappa_{1}\Gamma p_{11}+2p_{12}(\kappa_{2}-\pi),\\ \tilde{q}_{12}=\tilde{q}_{21}=\frac{1}{2}\kappa_{1}\Gamma p_{12}+p_{22}(\kappa_{2}-\pi)-\frac{1}{2}\Gamma p_{11},\\ \tilde{q}_{22}=-\Gamma p_{12}.\end{cases}$ (38)
Given that $\pi$ is bounded, $V_{2}$ is positive definite and $\dot{V}_{2}$ is negative definite for every value of the perturbation derivative $|\dot{\bm{\Pi}}|\leq\bm{\delta}$ subject to the following conditions
$\begin{split}&{p}_{11}=1,\ \ \ {p}_{22}>{p_{12}}^{2},\ \ \ {p}_{12}<0,\ \ \ \Gamma>0,\ \ \ \forall t\geq 0,\end{split}$ (39)
$\kappa_{1}\Gamma^{2}p_{11}p_{12}+\frac{1}{4}(\Gamma p_{11}-\kappa_{1}\Gamma p_{12})^{2}+2\Gamma p_{12}^{2}(\kappa_{2}-\pi)$
$-(p_{11}-\kappa_{1}p_{12})\Gamma p_{22}(\kappa_{2}-\pi)+(\kappa_{2}-\pi)^{2}p_{22}^{2}<0.$ (40)
Noting that $(\kappa_{2}-L)\leq(\kappa_{2}-\pi)\leq(\kappa_{2}+L)$, with $L$ defined in (13). When $L<\pi$, then $|s_{i}|<\mu_{\alpha}$ and from (13), $L$ will increase until the sliding mode occurs or $L>\pi$ is established. To maintain the sliding mode and ensure the boundedness around the origin for perturbation $\pi(t,y)=L$, $\kappa_{2}>L$ must be satisfied. Therefore, the condition (40) will be satisfied if
$\kappa_{1}\Gamma^{2}p_{12}+\frac{1}{4}(\Gamma-\kappa_{1}\Gamma p_{12})^{2}+2\Gamma p_{12}^{2}(\kappa_{2}+L)\\ -(1-\kappa_{1}p_{12})\Gamma p_{22}(\kappa_{2}-L)+(\kappa_{2}+L)^{2}p_{22}^{2}<0.$ (41)
Define
$\begin{cases}\gamma=\frac{p_{22}}{p_{12}^{2}},\ \ \theta=\frac{\kappa_{2}-L}{\kappa_{2}+L},\\ \Omega_{1}=-\kappa_{1}p_{12},\ \ \Omega_{2}=p_{22}(\kappa_{2}+L),\end{cases}$ (42)
and the condition (41) is satisfied if
[…]
is met. The first inequality in (43) represents the interior of a size-variable ellipse in the plane of $\Omega_{1}$ and $\Omega_{2}$, of which the size is determined by $\gamma$, $\theta$, and $\Gamma$. Functions $\Omega_{1},\Omega_{2}\in\mathbb{R}^{+}$ must be selected to satisfy (43), which requires $\gamma>1$, and is always possible if $\gamma\theta>1$.
Given the parameters $\theta,\gamma,\Omega_{1},\Omega_{2}$ that satisfy (43), the components of (34) that satisfy (39) and (40) are given as
${P}:\begin{cases}{p}_{11}=1,\\ {p}_{12}=p_{21}=-\sqrt{\frac{(1-\theta)\Omega_{2}}{2\gamma L}}=-\sqrt{\frac{p_{22}}{\gamma}},\\ {p}_{22}=\frac{(1-\theta)\Omega_{2}}{2L}.\end{cases}$ (44)
The variable gains $\kappa_{1}$ and $\kappa_{2}$, provided values of $\theta$ and $\gamma$, are given as
$\kappa_{1}=\Omega_{1}\sqrt{\frac{2\gamma}{(1-\theta)\Omega_{2}}}\sqrt{L},\ \ \kappa_{2}=\frac{\theta+1}{1-\theta}L,$ (45)
with $\Omega_{1}$ and $\Omega_{2}$ selected as any points within the ellipse (41), which exist in the first quadrant. This selection assures robust, finite-time boundedness about the origin of the super-twisting algorithm.
Case 1a: For the case when $\Gamma\neq 0$, implying $\dot{\bm{e}}\neq 0$, using (36) we have that
$\dot{V}_{2}\leq-|y_{1}|^{-\frac{1}{2}}\lambda_{min}(\tilde{Q}_{R})||\bm{\zeta}||_{2}^{2},$ (46)
where $\tilde{Q}_{R}\in\mathbb{R}^{+}$ is computed as
${\tilde{Q}_{R}}:\begin{cases}\tilde{q}_{R_{11}}=\kappa_{1}\Gamma-2\Omega_{2}\frac{p_{12}}{p_{22}}+(1+\Omega_{1})\Omega_{2}\frac{1-\theta}{p_{12}},\\ \tilde{q}_{R_{12}}=\tilde{q}_{R_{21}}=\Omega_{2}-\frac{1}{2}(\Gamma+\Gamma\Omega_{1}),\\ \tilde{q}_{R_{22}}=-\Gamma p_{12},\end{cases}$ (47)
and remains positive definite as long as (43) holds.
Using (35), (46) and the fact that $|y_{1}|^{\frac{1}{2}}\leq||\bm{\zeta}||_{2}\leq\frac{V_{2}^{\frac{1}{2}}}{\lambda^{\frac{1}{2}}_{\textrm{min}}(P)}$,
$\dot{V}_{2}$ $\leq-|y_{1}|^{-\frac{1}{2}}\frac{\lambda_{\textrm{min}}(\tilde{Q}_{R})}{\lambda_{\textrm{max}}(P)}V_{2}$
$\leq-\frac{\lambda^{\frac{1}{2}}_{\textrm{min}}(P)\lambda_{\textrm{min}}(\tilde{Q}_{R})}{\lambda_{\textrm{max}}(P)}V^{\frac{1}{2}}_{2}\leq-\chi V^{\frac{1}{2}}_{2},$ (48)
where $\chi=\frac{\lambda^{\frac{1}{2}}_{\textrm{min}}(P)\lambda_{\textrm{min}}(\tilde{Q}_{R})}{\lambda_{\textrm{max}}(P)}>0$. From [26, Lemma 2], it is possible to assume that there exists a positive constant $L^{*}$ such that $L\leq L^{*}$ and $\forall t>0$. Consider the Lyapunov function (32), taking the derivative along the trajectories of the perturbed system while using (48) and the gain $L_{i}$ in (13) gives
$\dot{V}_{1}$ $=\dot{V}_{2}+\dot{L}(L-L^{*})$ (49)
$\leq-\chi V^{\frac{1}{2}}_{2}-\underbrace{\omega_{\alpha}|s|\textrm{sign}(|s|-\mu_{\alpha})|L-L^{*}|}_{\begin{subarray}{c}\Upsilon\end{subarray}},$ (50)
Suppose that $|s|>\mu_{\alpha}$, this gives
$\dot{V}_{1}\leq-\chi V^{\frac{1}{2}}_{2}-\omega_{\alpha}\mu_{\alpha}|L-L^{*}|.$ (51)
Note that $\omega_{\alpha}>0$ and $\mu_{\alpha}>0$. Jensen’s inequality for two variables, with $a\geq 0$, $b\geq 0$, and $0<m<n$, implies that $(a^{n}+b^{n})^{\frac{1}{n}}\leq(a^{m}+b^{m})^{\frac{1}{m}}$. Setting $n=2$, where $m=1$
$\dot{V}_{1}\leq-(\chi^{2}V_{2}+2\omega_{\alpha}^{2}\mu_{\alpha}^{2}(\frac{1}{2}|L-L^{*}|^{2}))^{\frac{1}{2}}\leq-\vartheta V_{1}^{\frac{1}{2}},$ (52)
where $\vartheta=\textrm{min}\{\chi,\sqrt{2}\omega_{\alpha}\mu_{\alpha}\}$. Therefore, the finite time converge to the region $|s|\leq\mu_{\alpha}$ is guaranteed from any initial condition $|s(0)|>\mu_{\alpha}$. In this case, the states $\dot{e}$ will be driven onto the sliding manifold $s$ before the states $\dot{e}$ converge to the origin within a domain of $|s|\leq\mu_{\alpha}$. Using Lemma 1, the upper bound on the reaching time is given as
$t_{f}(y_{0})\leq\frac{2}{\vartheta}V_{1}^{\frac{1}{2}}(y_{0}),$ (53)
where $y_{0}$ is the initial value of $y$. Experimental evaluation has shown that a value of $\mu_{\alpha}$ can be selected slightly higher than the measurement noise level on the corresponding position or rotation measurement output ($\epsilon_{c}<\mu_{\alpha}$) to attain a dynamic reduction of controller gains.
Suppose that $|s|<\mu_{\alpha}$. In this scenario, the function $\Upsilon$ in (50) can be negative. Consequently, $\dot{V}_{1}$ would be sign-indefinite, and the closed-loop boundedness of the $s$-dynamics cannot be concluded. However, when $s$ escapes the boundary ($s>\mu_{\alpha}$), under (13), $L$ immediately begins to increase again to keep $s$ in the domain. Thus, $L$ continues changing its value, and $s$ remains around the boundary layer limit $|s|=\mu_{\alpha}$. In conclusion, $s$ initially converges to the domain $|s|<\mu_{\alpha}$ in finite time and subsequently remains in the region $|s|\leq\mu_{\alpha}$, verifying the sliding mode boundedness [27].
Case 1b: Consider the case when $\dot{e}=0$ and $s\neq 0$. From (43) and (66), as $\Gamma\rightarrow 0$ this vertically compresses the ellipse represented by (43) towards the $\Omega_{1}$ axis, shown in Fig. 2. This may result in an obstruction of the reachability of (10).
Figure 2: Ellipses expressing the stable boundary of (43) for different values of $\Gamma$. $A$ is not bounded for the given parameters, $B$ is bounded for $\Gamma=0.5$ and unbounded for $\Gamma=0.25$, and $C$ is bounded for both $\Gamma=0.5$ and $\Gamma=0.25$.
For the case when $\dot{e}=0$ and $s\neq 0$, (24) becomes
$\ddot{e}$ $=-\kappa_{1}|s|^{\frac{1}{2}}\textrm{sign}(s)+\nu+\Pi$
[…]
Since $\kappa_{2}>|\dot{\Pi}|$, ensuring boundedness around the origin, $\ddot{e}\neq 0$ indicates that $\dot{e}=0$ is not an attractor in the reaching phase. Hence, the reachability of the sliding manifold (10) to a boundary in the finite time as $t_{f}(y_{0})$ in (53) is still guaranteed. Therefore it will not interfere with the convergence of $V_{1}$.
III-C3 State Convergence Analysis
The following analyses consider the state tracking stability and convergence of $e$ and $\dot{e}$, first without noise and second with noise, when the absolute second order sliding mode (SOSM), $\dot{s}=s=0$, is not possible.
Case 2a: Consider the case with no measurement noise. From (10), we have
$e=s-\beta|\dot{e}|^{\alpha}\textrm{sign}(\dot{e}).$ (55)
When SOSM ($\dot{s}=s=0$) occurs in finite time in the absence of measurement noise, (55) reduces to
$e=-\beta|\dot{e}|^{\alpha}\textrm{sign}(\dot{e}).$ (56)
Considering the definition of $\alpha\in(1,2)$: $\textrm{diag}(|\dot{e}|^{\alpha})\textrm{sign}(\dot{e})=\dot{e}^{\alpha}$ for $\dot{e}>0$ and $\dot{e}<0$, the closed-loop dynamics is
$\dot{e}=-\beta^{-\frac{1}{\alpha}}(e)^{\frac{1}{\alpha}},$ (57)
and the finite time convergence of $e$ and $\dot{e}$ is achieved.
Consider the Lyapunov function candidate as
$V_{3}=\frac{1}{2}e^{2}.$ (58)
Taking the time derivative of (58) and substituting (57) gives
$\dot{V}_{3}$ $={e}\dot{e}={e}\left(-\beta^{-\frac{1}{\alpha}}(e)^{\frac{1}{\alpha}}\right)=-\beta^{-\frac{1}{\alpha}}(e)^{\frac{1+\alpha}{\alpha}}$
$\leq-\beta^{-\frac{1}{\alpha}}2^{\frac{1+\alpha}{2\alpha}}V_{3}^{\frac{1+\alpha}{2\alpha}}.$ (59)
Therefore, we obtain that $e$ and $\dot{e}$ can converge to the origin along the sliding surface in finite time as long as $\beta>0$ and $21<\alpha<$ 2. With $\alpha\in(1,2)$, this implies $\frac{\alpha+1}{2\alpha}\in(\frac{3}{4},1)$. Using Lemma 1, the upper bound on the convergence time is
$t_{f}(e)\leq\frac{2\alpha}{\beta^{-\frac{1}{\alpha}}(\alpha-1)}\frac{V_{3}(e(0))^{\frac{\alpha-1}{2\alpha}}}{2^{\frac{1+\alpha}{2\alpha}}}.$ (60)
Case 2b: In the presence of bounded Lebesgue-measurable noise, it was shown in Section III-C that the sliding variable $s$ will converge to a region $|s|\leq\mu_{\alpha}$, which is guaranteed from any initial condition $|s(0)|>\mu_{\alpha}$. With these considerations, from (10) we get
${e}+\beta|\dot{e}|^{\alpha}\textrm{sign}(\dot{e})=s,\ \ |s|\leq\mu_{\alpha},$ (61)
which can be rewritten, when $|\dot{e}|\neq 0$, as
${e}+\underbrace{\left(\beta-\frac{s}{|\dot{e}|^{\alpha}\textrm{sign}(\dot{e})}\right)}_{\begin{subarray}{c}\pi\end{subarray}}|\dot{e}|^{\alpha}\textrm{sign}(\dot{e})=0.$ (62)
When $\pi>0$, then Eq. (62) keeps the NTSM form considered in (56) and the finite time convergence is maintained and the error is bounded. Consequently, from $\pi$, the velocity tracking error converges to a region
$|\dot{e}|\leq\left(\frac{\mu_{\alpha}}{\beta}\right)^{\frac{1}{\alpha}}.$ (63)
Additionally, from (61)--(63), we can conclude that the position error will converge to a region
$|{e}|\leq\beta|\dot{e}|^{\alpha}+|s|\leq 2\mu_{\alpha},$ (64)
in finite time. This concludes the proof.
Remark 1
In practical applications, $\mu_{\alpha}$ is set based on acceptable error bounds, informed by factors like sampling intervals and system noise. If $\mu_{\alpha}$ is set too low relative to these factors, the error may not converge within the defined bounds, leading to a continuous increase of the adaptive gains. The parameter $\omega_{\alpha}$ requires careful calibration to balance quick adaptability to dynamic disturbances and noise against the risk of inducing undesirable oscillations. Other parameters such as $\eta_{\alpha}$, $\kappa_{1\text{max}}$, and $\kappa_{1\text{min}}$, although less critical, should be tuned to the specific application to ensure optimal performance.
Remark 2
The boundedness proof for the adaptive NT-STSM controller offers more flexibility in tuning compared to recent similar applications as in [9]. This proposed method allows for a range of acceptable ratios between $\kappa_{1}$ and $\kappa_{2}$ within defined regions $\Omega_{1}$ and $\Omega_{2}$, in contrast to the strict relationship required in [9], where $\kappa_{2}=\frac{1}{4}\Gamma\kappa_{1}$. The relation between $\kappa_{1}$ and $\kappa_{2}$ imposed by Theorem 1 is obtained by rearranging (45):
$\kappa_{2}=\frac{\Omega_{2}(\theta+1)}{2\gamma\Omega_{1}^{2}}\kappa_{1}^{2}.$ (65)
Conservatively, $\Omega_{1}$ and $\Omega_{2}$ can be selected as the center of the ellipse represented by (43), which is given as
$\Omega_{1c}=\frac{\gamma-2\theta+\gamma\theta^{2}}{\gamma(1-\theta^{2})},\ \ \Omega_{2c}=\Gamma\frac{\theta\gamma-1}{\gamma(1-\theta^{2})}.$ (66)
However, as illustrated in Fig. 2 and (66), as $\bm{\Gamma}$ decreases, the ellipse is compressed and the viable region for $\Omega_{1}$ and $\Omega_{2}$ shrinks. From this observation, choosing small values of $\Omega_{1}$ and $\Omega_{2}$, located in the bottom left region of the ellipse ensures stability as $\bm{\Gamma}$ decreases. While this is more restrictive than the traditional STSM control [2], in which any values of $\Omega_{1}$ and $\Omega_{2}$ within the ellipse can be chosen, it is less restrictive than the relationship in [9] and allows for the practical implementation of NT-STSM control on hardware.
IV Simulation and Experimental Results
IV-A Implementation
Simulations and experiments are performed using a Franka Emika 7-DOF manipulator and Robot Operating System (ROS). Simulations are conducted in Gazebo, based on the dynamics outlined in [28]. The controllers are run at a sampling rate of $\Delta t=0.001$ s. The dynamic parameters $\mathcal{M}$, $\mathcal{C}$, $\mathcal{G}$, and $J$ are computed internally by the manipulator software.
The proposed NT-STSM controller is compared with a proportional-derivative (PD) controller [29], NTSM controller [1], and STSM controller [3], formulated as
$\bm{u}_{\textrm{PD}}=$ $\ -K_{p}\bm{e}-2\sqrt{K_{p}}\dot{\bm{e}}+\bar{\mathcal{C}}+\bar{\mathcal{G}},$
$\bm{u}_{\textrm{NTSM}}=$$\ \bar{\mathcal{M}}\left(\ddot{\bm{x}}_{d}+H^{-1}\left(\frac{-\dot{\bm{e}}^{(2-\alpha)}}{\alpha\beta}-\kappa_{1}\textrm{sign}(\bm{s}_{1})\right)\right)$
$+\bar{\mathcal{C}}+\bar{\mathcal{G}},$
$\bm{u}_{\textrm{STSM}}=$$\ \bar{\mathcal{M}}\left(\ddot{\bm{x}}_{d}-H^{-1}\left(\kappa_{1}|S_{2}|^{\frac{1}{2}}\textrm{sign}(\bm{s}_{2})+\bm{\nu}\right)\right)+\bar{\mathcal{C}}+\bar{\mathcal{G}},$
$\dot{\bm{\nu}}=$ $\ -\kappa_{2}\textrm{sign}(\bm{s}_{2}),$
where $\bm{s}_{1}=\ \bm{e}+\beta\dot{\bm{e}}^{\alpha}$ and $\bm{s}_{2}=\ \bm{e}+\beta\dot{\bm{e}}$.
To smooth the control signal and mitigate the chattering in SM controllers, the $\textrm{sign}(\bm{s})$ terms are substituted with $\textrm{tanh}(k_{s}\bm{s})$, where $k_{s}$ is a positive design parameter that provides a trade-off between tracking accuracy and convergence speed. To ensure a fair comparison, all SM controllers were implemented using the same adaptive gain law (13) with the same parameters, as shown in Table I. Three sets of PD controller gains were tested to determine a fair baseline to compare with the proposed controller, shown in Table I.
TABLE I: Controller and observer parameters.
PD-low $K_{p\textrm{trans}}=200$ $K_{p\textrm{rot}}=50$
PD-med $K_{p\textrm{trans}}=800$ $K_{p\textrm{rot}}=200$
PD-high $K_{p\textrm{trans}}=2000$ $K_{p\textrm{rot}}=500$
Sliding-Mode $\beta=1$ $\alpha=9/7$ $k_{s}=30$
Parameters $\theta=0.9$ $\gamma=6$ $\Omega_{1}=1.5$
Eqs. (10), (14) $\Omega_{2}=0.14$
Adaptive $\kappa_{1\textrm{min}}=5$ $\kappa_{1\textrm{max}}=200$ $\omega_{\alpha}=1000$
Parameters $\mu_{\alpha}=0.001$ $\eta_{\alpha}=0.1$
Eq. (13)
Observer $\bar{F}=\begin{bmatrix}20,&20,&20,&20,&20,&20\end{bmatrix}^{T}$
Parameters $\alpha_{o0}=1.1\bar{F}$ $\alpha_{o1}=1.5\bar{F}^{\frac{1}{2}}$ $\alpha_{o2}=1.9\bar{F}^{\frac{1}{3}}$
Eqs. (5), (6), (7) $k_{o1}=200$ $k_{o2}=400$ $\alpha_{e}=0.02$
$\eta_{q}=0.5$
To evaluate the robustness of the controllers to disturbances, the simulations are conducted with significant joint friction, unknown to the controllers. The friction parameters are detailed in Table II, where $c$ is the damping value, $\mu_{s}$ is the static friction value, $\mu_{k}$ is the Coulomb friction coefficient, and $\mu_{v}$ is the viscous friction coefficient for each joint. Additionally, in both simulation and experiment, the dynamics of the manipulator are not exactly known.
Figure 3: Franka Emika 7-DOF manipulator. (a) Gazebo simulation environment; (b) Hardware setup grasping an object.
TABLE II: Simulation joint friction parameters.
[…]
Default 0.003 0 0 16
Simulated 0.003 0.5 25 25
In the simulation, the robot executes a simultaneous motion of $0.05$ m translation in the $x$, $y$, and $z$ axes and $25^{\circ}$ rotation about the $x$, $y$, and $z$ axes. In the experiment, the robot manipulates a time-varying payload of a container holding steel balls with a total mass of 0.5 kg. The robot pours out the contents by executing a $0.1$ m translation in the $y$-axis and a $40^{\circ}$ rotation about the $z$-axis. Fig. 3 shows the manipulator in the simulation and experimental setups.
IV-B Simulation Results
Fig. 4 compares different velocity estimation methods. The Franka robot’s built-in velocity measurements (Act) are derived from position data, making them inherently noisy. Applying a low-pass filter (Smooth) reduces noise but introduces a measurement delay. During testing, both noise and delay in velocity estimates led to chattering in the task-space sliding-mode controllers. To mitigate this, the TOSM observer was implemented, providing a balance between accuracy, noise reduction, and minimal delay. Fig. 4 also highlights why applying SM control is more challenging for rotational motion. From 0-5 s, when only translational motion occurs, the velocity noise is low. However, during rotational motion (9-13 s and 16-20 s), the noise is significantly higher - even in translational velocity, increasing susceptibility to chattering.
Figure 4: Comparison of x-axis velocity measurements.
Figs. 5 and 6 show the position and orientation tracking results from the simulation. As expected, increasing the PD gains improved the tracking accuracy but increased the control effort, shown in Table III, where the performance metrics are Root Mean Squared Error (RMSE) and average control effort, $\tau_{\textrm{avg}}$. Note that the PD-low controller is significantly affected by the frictional disturbances, resulting in poor tracking.
Despite chattering reduction techniques, the NTSM controller exhibited severe chattering and is omitted. The STSM controller performed well until 14 s, when it began to chatter significantly until stabilizing at 17 s. In contrast, the proposed NT-STSM controller, using the same parameters, achieved more accurate translational and rotational tracking comparable to the PD-high controller while maintaining a reasonable level of control effort. The NTSM and STSM could be tuned to reduce chattering, however, this would require reducing $\kappa_{1}$ and $\kappa_{2}$, which would reduce the tracking accuracy and not be comparable to the accuracy of the NT-STSM controller.
Applying the NT-STSM controller from [9], originally designed for the joint space, to the task space with its constrained relationship between $\kappa_{1}$ and $\kappa_{2}$, resulted in a 71.3% increase in $\textrm{RMSE}_{p}$, and a 71.1% increase in $\textrm{RMSE}_{\xi}$, compared to our approach. Compared to the PD-med controller, the proposed NT-STSM controller reduced $\textrm{RMSE}_{p}$ by 29.0% and $\textrm{RMSE}_{\xi}$ 30.4%, with similar control effort. Compared to the PD-high controller, the proposed NT-STSM controller achieved similar tracking performance with 13.4% lower control effort.
Figure 5: Simulation on the translational position tracking.
Figure 6: Simulation on the rotational position tracking.
TABLE III: Performance metrics for simulation tracking task.
Controller $\textrm{RMSE}_{p}$ $\textrm{RMSE}_{\xi}$ $\tau_{\textrm{avg}}$
[m] [unitless] [Nm]
PD-low 7.25e-3 1.16e-2 3.69e-1
PD-med 2.48e-3 7.34e-3 5.06e-1
PD-high 1.54e-3 5.88e-3 5.66e-1
NTSM N/A N/A N/A
STSM 7.04e-3 2.71e-2 1.92
NT-STSM [9] 6.13e-3 1.77e-2 5.10e-1
NT-STSM (ours) 1.76e-3 5.11e-3 4.90e-1
To further demonstrate the robustness of the NT-STSM controller, the same task was performed under a series of external disturbances applied to the end-effector. At $5$ s, $\bm{f}_{e}=\begin{bmatrix}0&5&0&0&0&0\end{bmatrix}$, at $10$ s, $\bm{f}_{e}=\begin{bmatrix}5&0&-5&0&0&0\end{bmatrix}$, at $14$ s, $\bm{f}_{e}=\begin{bmatrix}0&0&0&1&0&0\end{bmatrix}$, and at $17$ s, $\bm{f}_{e}=\begin{bmatrix}0&-5&5&0&-1&1\end{bmatrix}$, where the first three elements of $\bm{f}_{e}$ have units of N and the last three have units of Nm. The disturbances are each applied for 1 s from each starting time. Figure 7 compares translational tracking performance, and Table IV shows the performance metrics, demonstrating that the PD-med controller is more affected by disturbances despite having similar control effort as the NT-STSM controller.
Figure 7: Simulation on the translational position tracking with external disturbances. The red dashed lines indicate when external forces are applied.
TABLE IV: Performance metrics for simulation tracking task with external disturbances.
Controller $\textrm{RMSE}_{p}$ $\textrm{RMSE}_{\xi}$ $\tau_{\textrm{avg}}$
[m] [unitless] [Nm]
PD-med 2.72e-3 7.23e-3 7.13e-1
NT-STSM (ours) 1.89e-3 6.16e-3 7.16e-1
IV-C Experimental Results
A video of the experiment is available at https://www.youtube.com/watch?v=ziweWAJxdYU. Fig. 8 and Table V compare the tracking performance of the PD-med and the NT-STSM controllers along the $y$ and $z$ translational axes and the $\epsilon_{z}$ rotational component. While the PD controller achieves lower tracking error in the $y$ direction and $\epsilon_{z}$ orientation, the NT-STSM controller demonstrates significantly improved tracking along the $z$-axis, which is most sensitive to the payload. This is clear at $t=18$ s, where the payload change degrades the PD controller’s performance while the NT-STSM controller maintains more accurate tracking. Additionally, the control input of the NT-STSM is significantly lower than the PD-med. These results confirm the robustness and efficiency of the proposed NT-STSM controller in the presence of unknown disturbances or varying payload and validate its viability for practical robotic applications involving external perturbations.
Figure 8: Experimental pouring task with selected position and orientation tracking. The payload starts changing at $t=18$ s.
TABLE V: Performance metrics for experimental pouring task.
Controller $\textrm{RMSE}_{py}$ $\textrm{RMSE}_{pz}$ $\textrm{RMSE}_{\xi z}$ $\tau_{\textrm{avg}}$
[m] [m] [unitless] [Nm]
PD-med 7.30e-3 7.23e-3 1.36e-2 1.30
NT-STSM (ours) 9.41e-3 2.74e-3 1.42e-2 8.83e-1
V Conclusion
This paper proposed an adaptive NT-STSM controller in task space for a 7-DOF robotic manipulator. A novel boundedness analysis established that tracking errors converge within a known region, providing flexible tuning guidelines. Compared to conventional controllers, the proposed method improves robustness and mitigates chattering, validating its application for real-time task-space control. Future work will incorporate force feedback and apply the controller to multi-manipulator systems based on the convincing results of this work.